A price p

 (in dollars) and demand x

 for a product are related by

2x^2-8xp+50p^2=7400



If the price is increasing at a rate of 2 dollars per month when the price is 10 dollars, find the rate of change of the demand.


1
Expert's answer
2021-11-15T16:40:26-0500

"\displaystyle
2x^2 -8xp+2x^2=7400, \frac{dp}{dt} = 2\\
\text{Differentiating implicitly with respect to t, we have}\\
\frac{d}{dt}(2x^2-8xp+50p^2) = \frac{d}{dt}7400\\
4x\frac{dx}{dt}- 8p\frac{dx}{dt}- 8x\frac{dp}{dt}+100p\frac{dp}{dt}=0-(1)\\
\text{Recall $\frac{dp}{dt}=2$}\\
\text{Therefore 1 becomes}\\
\frac{dx}{dt}= \frac{16x-200p}{4x-8p}-(2)\\
\text{Next, we want to obtain the values of x and p, at p = 10}\\
2x^2-80x+5000=7400\\
=2x^2-80x-2400=0\\
\implies \text{x = 60 and x = -20}\\
\text{The only feasible value for x is 60, since x is demand and can't be }\\
\text{negative}\\
\text{Hence p = 10} \implies x =60\\
\text{Substituting in (2), we have that}\\
\frac{dx}{dt}= \frac{16(60)-200(10)}{4(60)-8(10)}=-6.5\\
\text{Therefore, the demand decreases with time at rate of 6.5}"